Operations Report – January 5th

Crew 272 Operations Report 5-1-2023

SOL: 4

Name of person filing report: Mason Kuhn

Non-nominal systems: None!

Notes on non-nominal systems: The previously mentioned Hab toilet noise does not seem to be an issue anymore.


Spirit rover used: No

Hours: 206.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 110.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 218.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: None were used today, no EVA

Summary of Hab operations: I gathered the upper deck trash and toilet room trash. It has been placed in the rear airlock for pickup.

WATER USE: 19.6 gallons

Water (static tank): 311.5 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: All suits 1-11 are currently charging. All radios are charged and off of the charging docks.

Summary of GreenHab operations: Watered and monitored all crops. Harvested herbs for use in tonight’s ravioli dinner, beans and microgreens for snacks, and dill to include in tomorrow’s bread. Watered experimental crops, and watered new crops in regolith and soil experiments. Planted black nebula carrots.

WATER USE: 7 gallons total

Heater: On

Supplemental light: 4 hours

Harvest: 10g microgreens, 5g thyme, 2g basil, 3g chive, 15g beans, 25g dill

Summary of ScienceDome operations: The Crew Geologist investigated many rock & mineral samples from our previous EVAs, and several other crew members joined to watch and learn from her. The ring light magnifier in the SciDome was used.

Dual split: On

Summary of RAM operations: Retrieved WD-40 to use on a stateroom’s squeaky door hinge, then put the WD-40 back in the RAM. Yesterday evening, I recognized a low-battery chirp from a CO alarm in the RAM (1 chirp, once per minute). Today I went to replace the 9V in it, but it was not chirping anymore. I will continue to check every sol to see if the chirping returns.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Request for trash to be taken from the rear airlock in the morning
  • Request for information on the usability of the sink in the ScienceDome


Mason Kuhn

Crew Engineer


Copyright © The Mars Society. All rights reserved. | Main Site